OpenStudy (anonymous):

Find the derivative of f(x) = sin^-1(x/2). Leave your answer in simplified form.

OpenStudy (anonymous):

f'(x)= -1/2 [cos^(-3x/2)] f'(x)=-1/[2 sec^(3x/2)]

myininaya (myininaya):

\[y=\sin^{-1}(\frac{x}{2})=>\sin(y)=\frac{x}{2}\] take derivative of both sides \[\cos(y) y'=\frac{1}{2}=>y'=\frac{1}{2 \cos(y)}\] but sin(y)=opp/hyp=x/2 using pythagorean thm we can find the adj side of y so we have adj=sqrt{2^2-x^2}=sqrt{4-x^2} so \[y'=\frac{1}{2 \frac{\sqrt{4-x^2}}{2}}=\frac{1}{\sqrt{4-x^2}}\]